Schedule of Cash Collections of Accounts Receivable

OfficeMart Inc. has "cash and carry" customers and credit customers. OfficeMart estimates that 25% of monthly sales are to cash customers, while the remaining sales are to credit customers. Of the credit customers, 25% pay their accounts in the month of sale, while the remaining 75% pay their accounts in the month following the month of sale. Projected sales for the next three months are as follows:

October $107,000
November 134,000
December 196,000

The Accounts Receivable balance on September 30 was $72,000.

Prepare a schedule of cash collections from sales for October, November, and December. Round all calculations to the nearest whole dollar.

OfficeMart Inc.
Schedule of Cash Collections from Sales
For the Three Months Ending December 31
October November December
Receipts from cash sales:
Cash sales $ $ $
September sales on account:
Collected in October
October sales on account:
Collected in October
Collected in November
November sales on account:
Collected in November
Collected in December
December sales on account:
Collected in December
Total cash receipts $ $ $

please show work
Answer #1

Answers

OfficeMart Inc.

Schedule of Cash Collections from Sales

For the Three Months Ending December 31

October

November

December

Receipts from cash sales:

Cash sales

$26,750

$33,500

$49,000

September sales on account:

Collected in October

$72,000

October sales on account:

Collected in October

$20,063

Collected in November

$60,188

November sales on account:

Collected in November

$25,125

Collected in December

$75,375

December sales on account:

Collected in December

$36,750

Total cash receipts

$118,813

$118,813

$161,125

--Working

OfficeMart Inc.

Schedule of Cash Collections from Sales

For the Three Months Ending December 31

October

November

December

Receipts from cash sales:

Cash sales

=107000*25%

=134000*25%

=196000*25%

September sales on account:

Collected in October

72000

October sales on account:

Collected in October

=107000*75%*25%

Collected in November

=107000*75%*75%

November sales on account:

Collected in November

=134000*75%*25%

Collected in December

=134000*75%*75%

December sales on account:

Collected in December

=196000*75%*25%